OpenStudy (anonymous):

A QUADRATIC EQUATION QUESTION The total resources T (in billions of dollars) of the Pension Benefit Guaranty Corporation, the government agency that insures pensions can be approximated by the equation: T=-0.26x^2+3.62x+30.18 Where x is the number of years after 2000. Determine when the total resources are at the given level. 1.$30 billion 2.When will the Corporation be out of money (T=0)?

OpenStudy (whpalmer4):

for part 1, solve \[30 = -0.26x^2+3.62x+30.18 \] for \(x\) For part 2, solve \[0 = -0.26x^2+3.62x + 30.18\] for \(x\) Quadratic formula is probably the way to go: \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]given \[ax^2+bx+c=0, a\ne 0\]

OpenStudy (anonymous):

please can you slove the part 2 for me.

OpenStudy (whpalmer4):

in part 2, we have \(a=-0.26, b = 3.62, c = 30.18\) so our solutions are \[x=\frac{ -3.62\pm\sqrt{3.62^2-4(-0.26)(30.18)}}{2*(-0.26)} =\]

OpenStudy (anonymous):

oh ok.got you now you so good.can i ask you for another help with another question?

OpenStudy (whpalmer4):

what's the question?

OpenStudy (anonymous):

A well-known golfer sells his book Help of Hackers for $13.25 per copy. His fixed costs are $125,000 and he estimates that the variable cost of printing, binding, and distribution are given by -4x^2+2880x-295. Find expressions for the revenue, cost, and profit from selling x thousand copies of the book.

OpenStudy (whpalmer4):

revenue is sales, so revenue = price per copy * x * 1000 (because x is in thousands of copies) cost is fixed cost + variable cost, and those are both given to you profit is revenue - costs, so just combine the two previous equations accordingly

OpenStudy (anonymous):

so u mean the revenue = $13.25x(thousand)

OpenStudy (whpalmer4):

The price per book is $13.25, but x=1 is 1000 books, so the revenue from x=1 would be $13250.

OpenStudy (whpalmer4):

so revenue = 13250x

OpenStudy (anonymous):

so can u please write out the cost for me.?want to see if im right

OpenStudy (whpalmer4):

you write out what you think it is, and I'll tell you if you have it right.

OpenStudy (anonymous):

$125000-4x^2+2880x-295

OpenStudy (whpalmer4):

Yes.

OpenStudy (anonymous):

so profit =13250x-125000-4x^2+2880x-295 or ?? i guess im wrong

OpenStudy (whpalmer4):

well, regardless of any question of whether you've added/subtracted the right pieces, you did the algebra wrong... \[revenue = 13250x\] \[cost = 125000-4x^2+2880x-295\] \[profit = revenue - cost = 13250x - (125000-4x^2+2880x-295)\]\[=13250x-125000-(-4x^2)-2880x-(-295)\]

OpenStudy (whpalmer4):

when you subtract a polynomial, you have to put parentheses around the whole thing and distribute that -1 factor...

OpenStudy (whpalmer4):

\[x+a - x -b = a-b\]\[x+a - (x-b) = a+b\]

OpenStudy (anonymous):

so after this,need i group like terms and subtract or add them accordingly or i should just leave the answer like that?

OpenStudy (whpalmer4):

oh, I would collect like terms and simplify, definitely!

OpenStudy (whpalmer4):

just make sure you don't make any mistakes in the process :-)

OpenStudy (whpalmer4):

Here's a graph of cost, revenue and profit. Can you figure out which line is which? How many thousand books need to be sold (x=) before there is a profit?

OpenStudy (anonymous):

wow you are soo helpful.

OpenStudy (whpalmer4):

Note the fact that the cost equation is a parabola makes pretty much no difference at all, because the constant factors are >> the value of x^2 for the range of interest.

OpenStudy (whpalmer4):

it's only when sales get out into the hundreds of thousands of copies that you can see it on the graph:

OpenStudy (whpalmer4):

were you able to figure out which line is which, and estimate the break-even point for selling enough copies to start making a profit?

OpenStudy (anonymous):

so after simplifying i need to make sure my x^2 comes first followed by my x number and then the number without any variable.just like how it appears in the variable cost?

OpenStudy (whpalmer4):

well, I don't think you have to write it in that order, but usually people do (or in the reverse order).

OpenStudy (anonymous):

oh ok.that would therefore be my final answer right after i have simplified right?

OpenStudy (whpalmer4):

Yes. Though to be honest, even without simplification or rearrangement you've satisfied the request to "Find expressions for the revenue, cost, and profit from selling x thousand copies of the book."

OpenStudy (anonymous):

oh kk.that would be all then .thank you very much you far too kind.im very grateful :-)

OpenStudy (whpalmer4):

If they want fully simplified expressions written in descending order (or the problem is marked incorrect), the problem statement should say so :-)

OpenStudy (whpalmer4):

You're welcome. I like to show people that there are actually interesting things going on in these problems besides just finding the solution :-)

OpenStudy (anonymous):

:-)

OpenStudy (whpalmer4):

For example, you can see in that cost graph that it costs a bit less to make each additional book once you're making a sufficient volume of them. volume discounts, baby!

OpenStudy (anonymous):

hahhahahahaha.i gotcha